OpenStudy (anonymous):

Multiply (a+g)(a^2-ag+g^2)

OpenStudy (anonymous):

\[a^3+g^3\]

OpenStudy (anonymous):

if you recognize it as the formula for factoring the sum of two cubes. otherwise you have to do 6 multiplications and a raft of cancellation

OpenStudy (anonymous):

(a+g)(a^2-ag+g^2) =a(a^2-ag+g^2)+g(a^2-ag+g^2) =a^3-a^2g+ag^2+a^2g-ag^2+g^3 =(a^3+g^3)